40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 2 19:59:54 2012, Jenne, Update, Alignment, POP, POX, POY, IPPOS, IPANG, REFL all coming out of vac 
    Reply  Fri Nov 2 20:41:53 2012, Jenne, Update, Alignment, AS, REFL camera shots AS_2Nov2012.pngREFL_2Nov2012.png
       Reply  Fri Nov 2 21:40:04 2012, Manasa, Update, Alignment, AS, REFL camera shots 
       Reply  Mon Nov 5 09:53:35 2012, Jenne, Update, Alignment, AS, REFL camera shots AS_5Nov2012.pngREFL_5Nov2012.pngIPPOS_IPANG_weekend_trend.png
    Reply  Sat Nov 3 10:31:59 2012, rana, Update, Alignment, POP, POX, POY, IPPOS, IPANG, REFL all coming out of vac 
Message ID: 7666     Entry time: Fri Nov 2 21:40:04 2012     In reply to: 7665
Author: Manasa 
Type: Update 
Category: Alignment 
Subject: AS, REFL camera shots 

 

 To get the camera shot of AS, Y1 mirror on the path was replaced by a 99% BS and transmitted beam was directed to the camera via a 50-50 BS (ND filters were distorting the image on the camera introducing fringes).

ELOG V3.1.3-